answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

If a homozygous black guinea pig is crossed with a white guinea pig what is the probability that an offspring will have black fur?

depends if the black fur gene is dominant..

if it is...

then yu would cross BB with bb making all heterozygous genotypes(Bb)

therefore, having all possible offspring with black fur

so theres a 100% probability of offspring with black fur(:

What are the odds of rolling a 7 at a craps table?

1 in 6 The way you calculate this is as follows: There are 36 possible combinations you can get with two dice. Six of those combinations add up to 7 and 30 of those combinations add up to either 6, 5, 4, 3, or 2. Therefore, there are 30 ways you can roll a number other than seven and six ways you can roll a seven. If you want a seven to come out, there are 30 ways to lose and 6 ways to win. This translates into odds of 30 to 6 or 5 to 1. The difference between this and the above answer is the way odds are stated. They can be stated either using "in" or "to." You can say the odds are either 6 ways to win "in" 36 possible combinations or you can say there are 30 ways to lose "to" 6 ways to win. When these are reduced, they become 1 way to win "in" 6 ways to lose and 5 ways to lose "to" 1 way to win. Thus, you get "1 in 6" which is equivalent to "5 to 1".

Four cards are drawn at random from a pack of 52 playing cards Find the probability of getting all face cards?

Face cards in a regular set of playing cards consist of Jack, Queen, King, and Ace. With these 4 face cards occurring in all 4 suits (hearts, diamonds, clubs, and spades), that gives 16 total face cards.

If we draw 4 cards without putting them back, the probabilities are calculated on each draw as (number of face cards remaining)/(total number of cards remaining)

1st draw: 16/52 = 0.307692308

2nd draw: 15/51 = 0.294117647

3rd draw: 14/50 = .28

4th draw: 13/49 = 0.265306122

These are the individual probabilities of each step independent of each other. To calculate the probability when depending on the success of the previous step, we simply multiply each step's probabilities together.

Probability of drawing N face cards in a row:

1: (16/52) = 0.307692308

2: (16/52)*(15/51) = 0.0904977376

3: (16/52)*(15/51)*(14/50) = 0.0253393665

4: (16/52)*(15/51)*(14/50)*(13/49) = 0.00672268908

This translates to approximately a 7 in 1000 chance of drawing 4 face cards in a row out of a deck of 52 cards!

How many different bridge hands of 13 cards can be dealt from a 52 card deck?

The entire deck is dealt out for each hand of bridge. The maximum number of different hands that can be dealt is 52 x 51x 50 x 49 x 48 x 47 x 46 x 45 x 44 x 43 x 42 x 41 x 40. This equals

3,954,242,643,911,240,000,000 possible different deals.

or

three-sextillion, nine-hundred-fifty-four-quintillion, two-hundred-forty-two-quadrillion, six-hundred-forty-three-trillion,nine-hundred-eleven-billion, two-hundred-forty-million possible bridge hands. (U.S and modern British)

or

three-thousand-nine-hundred-fifty-four-trillion, two-hundred-forty-two-thousand-six-hundred-forty-three-billion, nine-hundred-eleven-thousand-two-hundred-forty-million possible bridge hands. (Continental Europe and/or Traditional British)

Quite a few possible hands.

What is the probability of rolling a 1 or a 5 with 2 dice?

The odds are: 1 of 36

1-1, 1-2, 1-3, 1-4, 1-5, 1-6, 2-1, 2-2, 2-3, 2-4, 2-5, 2-6, 3-1, 3-2, 3-3, 3-4, 3-5, 3-6, 4-1, 4-2, 4-3, 4-4, 4-5, 4-6, 5-1, 5-2, 5-3, 5-4, 5-5, 5-6, 6-1, 6-2, 6-3, 6-4, 6-5, 6-6

Each die is independent from the other (what you roll with one die does not affect what you can roll with the other). This means that each die can land on any number from 1-6. The sample set CAN involve duplicates, meaning a 1-2 is a DIFFERENT result than 2-1.

Since each die is independent from the other, we calculate the probability of each event happening, then multiple these probabilities by each other. The probability of each event is 1/6, or 0.167, so the changes of rolling double six is 0.0279 or 3%

If you draw a card from a standard deck of cards what is the probability of drawing a ace and a jack?

The probability of drawing an ace from a standard deck of cards.

There are 52 cards in the deck (after discarding Jokers) and there are 4 aces.

So: 4 out of 52 makes the probability 1 in 13 (one in thirteen).

What are the chances of picking a black coloured card out of a pack of cards?

there is a 50% probability that you will pick a black card out of a full pack of playing cards

Which card is face card?

A card that has a face on it (King, Queen, Jack).

What is the probability of picking a 2 or 4 in a deck of 52 cards?

About 15%

Since there are (4) 2's and (4) 4's in a deck, the question is really, "what is the probability of picking one of 8 specific cards from a deck of 52?"

The answer is 8/52.

A card is drawn from a deck what is the probability that it is a nine or a club?

The answer i got was 1 3/26

The cards are red or black, so the chance of getting black is 1/2

The cards less than nine is 1-8. There are 4 of each value, so 8 cards times 4 suits equals 32. The chance of getting a card less than nine is 32/52, or 8/13

1/2 plus 8/13 is 13/26 plus 16/ 26, which equals 29/26, or 1 3/26

That is how i solved it.

If you are dealt 4 cards from a shuffled deck of 52 cards what is the probability that all 4 cards are clubs?

I am assuming that you are not replacing the cards in the deck after each draw.

The probability of drawing no clubs = (52-13)/52*(51-13)/51*(50-13)/50*(49-13)/49

p= 39/52*38/51*37/50*36/49 ~= 0.3038 = 30.38%

The probability of drawing at least one club = 1-0.3038=.6962 = 69.62%

Whats Ace of spades is in the deck?

The Ace of Spades is commonly thought of as the highest-ranking card in the deck of playing cards, although the actual value of the card varies from game to game. In popular myth and folklore, it is also known as "The Death Card."

How many queens ace or kings are there in a deck of cards?

There is 4 kings and 4 queens in a deck. So 4+4=8 kings and queens

What is the probability of drawing at least one ace from a poker deck on two draws if the first card is replaced before the second card is drawn?

The probability of drawing a red card and a spade in two cards is the probability of drawing a red card multiplied by the probability of drawing a spade, multiplied by 2 (as it doesn't matter which way around they are drawn).

The probability of drawing a spade is 1/13 as there are 4 spades and 52 cards.

The probability of drawing a red card after this is 26/51 if the spade was black, and 25/51 if the spade was red. This averages at 51/102

Multiply these probabilities together and then multiply by two and we get 51/663 which can be simplified to 1/13

How many different kinds of cards are in a deck?

There are 13 different face values, 4 of each.

There are 4 different suits, 13 of each.

Each card is a unique combination of face value and suit.

Also, there are face cards (jack, queen, king, and sometimes ace), and number cards, and sometimes aces are considered by themselves.

How many number 7 cards are in a deck of 52 cards?

There is only one of each individual card in a deck. While there are 13 cards of the heart suit, and 4 cards that have a value of seven out of the entire deck, only ONE card is a seven of hearts.

How many possible rolls are there in Yahtzee?

Each die is going to land on 1 of 6 faces, therefore, one die = 1/6

There are five dice in yahtzee so you multiply the six chances of the first die by the six chances of the second, by the six chances of the fourth, etc.

1/6 * 1/6 * 1/6 * 1/6 * 1/6 = 1/7776 When rolling five die, all at the same time, there are 7776 combinations of numbers you can get.

What does cutting a deck of cards mean?

Usually when you are playing cards, the person who is dealing will offer the person to his right a chance to cut the cards, meaning to re-order them by taking a portion of them off the top of the deck and placing them on the bottom. This is a tradition related to the prevention of cheating. Some people are very good at stacking cards, and the cut is offered to make sure that things are fair. It isn't always as polite as "would you prefer to cut the cards?" Sometimes it is just "wanna cut?" Either way, the offer is being made for you to re-order the cards. You can refuse and let the deal stand as is, or you can remove the top of the deck (however many cards you want) and place it on the table. That's the cut, and the dealer will then make what you removed the bottom of the stack, and proceed to deal. If you would like further information, there is an article about cutting cards here:http://en.wikipedia.org/wiki/Cut_(playing_cards)

What is the probability that a card drawn at random is a queen or a heart?

P(at least one heart out of two cards drawn) = 1 - P(neither being a heart)

= 1 - (39/52)(38/51) = 1 - (2 * 3 * 13 * 19)/(2 * 2 * 3 * 13 * 17) = 1 - 19/34

= (34 - 19)/34 = 15/34

P(at least one heart out of two cards drawn) = 15/34

Trending Questions
Picking cards from a deck is dependent or independent? What is sampling examine well known methods of probability and nonprobability sampling in statistic? What operation is used when calculating the probability that 2 events will happen and Find the theoretical probability of rolling a total of 7 as a sum of two dices? In four successive tosses a fair coin lands heads up each time What is likely to happen when the coin is tossed a fifth time? How do you find the probability of a given b? What is the probability of getting a sum of 2? Can two mutually exclusive events occur at the same time? How many ways can you arrange the letters in the word RACECAR? What is the probability of not drawing a black six from a deck of cards? How many three digits can be formed from the digits 0 1 2 3 4 5 and 6 if each digits can be used only once? True or false When most materials cool they get a little smaller? What is the probability of rolling a six sided die 20 times and getting a 5? What is the probability of flipping a coin four times and getting the same side all four times? What are numbers that are relatively prime to each other? A coin is tossed 3times .determine P(EF)whereEHESD ON 3rd toss .Fhead occur on 1st two toss? How many times can 8 go into 152? In a word aspiration what is the probability for selecting one letter as vowel? In a study using 9 samples and in which the population variance is unknown the distribution that should be used to calculate confidence intervals is? When did Tunguska event happen? How do you determine probability?